Suppose marginal utility of good X is 20 while its price is Rs. 4 per unit and marginal utility of Y good is 50 while its price is Rs.5 per unit .The individual to whom this information applies is spending 20 on each good .Is he maximizing his satisfaction.?

1
Expert's answer
2021-12-03T12:01:18-0500

We have the following information-

  • Marginal utility of good X: MUx = 20
  • Marginal utility of good y: MUy = 50
  • Px = Rs 4 per unit
  • Py = Rs 5 per unit
  • I = 20 on each good.

The first-order condition for maximizing utility is given as-

"MRS =\frac{Px}{Py}"

Where "MRS =\frac{MUx}{MUy}"

so

"\frac{MUx}{MUy} =\frac{ Px}{Py}\\=\frac{20}{50}=\frac{2}{5}\\\frac{Px}{Py}=\frac{4}{5}"

Since, MRS "≠\frac{Px}{Py}" So, the level of satisfaction is not maximized.